# Winding quotients for virtual period maps of rank 1

### v5

@@ Abstract.
We illustrate a rank 1 model of virtual period maps and their associated winding quotients. The winding quotient is a new notion that appears in the recent study of virtual period maps. It requires a reformulation of the classical inversion problem for the classical period maps. We answer to the new inversion problem by introducing a $q$-multiplicatively periodic function, whose pull-back to the winding covering space is the Weierstrass $\wp$-function up to a correction by Eisenstein series $E_{2}$. This function appears as the propagator on elliptic curves in mathematical physics.
@@ 1. Introduction
The classical theory of elliptic integral of the first kind for a Weierstrass family of elliptic curves $X\to S$ (e.g. [Si] Chap.1) defines a period map, which is a multi-valent map from the complement of the discriminant divisor $S\setminus D$ of the base space $S$ of the family to its range domain (so called, period domain). The range is isomorphic to the monodromy covering of $S\setminus D$, i.e. the quotient space of the universal covering $(S\setminus D)^{\sim}$ divided by the kernel of the monodromy representation of $\pi_{1}(S\setminus D,*)$. Then, the Jacobi’s inversion problem for the period map ([ibid] 1.13.) is to ask for a description of the inverse map from the range to the defining domain $S\setminus D$, and its answer were given by Eisenstein series (see [ibid] Section13)
@@
Actually, the pull-back of the function $\mathcal{W}(w,q)$ to the winding covering space has been studied in quite different contexts in mathematical physics either as the propagator $P(z)$ by Dijkgraaf [D] (4.44) on elliptic curves and the propagator $P_{0}^{\infty}(w_{1},w_{2};\tau,\overline{\tau})$ of BCOV theory by Si Li ([L1] Lemma 3.1, [L2] section 4 and [LZ] 3.3), or as the genus-one Bergman kernel function $B(z_{1},z_{2})$ in the topological recursion by Eynard [E] p.339 Examples (c.f. also Tyurin [Ty] p.185 Example. $\omega_{\alpha}$ and Fay [F] (46)). We do not know yet whether such connection of the winding phenomenon with mathmatical physics is common for non-geometric virtual period maps in general or it remains only in these test examples. We need to test further cases [S3, G-S].
@@ 2. The virtual period map $z^{\tau}$ on $\mathbb{C}_{z}^{\times}$
from the punctured $z$-plane $\mathbb{C}_{z}^{\times}$ as the domain of the map to the punctured plane $\mathbb{C}_{w}^{\times}$ as the range of the map. Obviously it is not a univalent map. So we define the map geometrically as follows. We first choose $1\in\mathbb{C}_{z}^{\times}$ as the base point in the domain, and define the value of the map $z^{\tau}$ at $1\in\mathbb{C}_{z}^{\times}$ to be 1. Then, the values of the map are defined by analytic continuations of $\exp\!\big(\tau\log(z)\big)$ along paths on the $z$-plane $\mathbb{C}_{z}^{\times}$ starting at 1. The monodromy of this map by the counter clock-wise turn around 0 of the $z$-variable is given by the multiplication of $\exp{\big(2\pi\sqrt{-1}\tau\big)}$ on $z^{\tau}$, i.e. its multiplication in the $w$-plane $\mathbb{C}_{w}^{\times}$. Since $|\exp{\big(2\pi\sqrt{-1}\tau\big)}|\!<\!1$, the monodromy group $\big(\exp{(2\pi\sqrt{-1}\tau)}\big)^{\mathbb{Z}}$ is an infinite cyclic group and gives a faithful representation of the fundamental group $\pi_{1}(\mathbb{C}_{z}^{\times},1)\simeq\mathbb{Z}$. Then the monodromy covering space of the defining domain of the map $z^{\tau}$ is the universal covering space $\mathbb{C}_{\mathrm{Log}z}$ of $\mathbb{C}_{z}^{\times}$, where we denote by $\mathrm{Log}z$ the coordinate of the universal covering space such that $\exp(\mathrm{Log}z)=z$. On the space $\mathbb{C}_{\rm{Log}z}$, we have the covering transformation group action
@@ 3. Lifting of the map $z^{\tau}$ to the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$
Recall that $\mathrm{Log}z$ is the coordinate of the monodromy covering plane of $\mathbb{C}_{z}^{\times}$. Then, the virtual period map $z^{\tau}$ is lifted as the univalent map $w=\exp\big(\tau\mathrm{Log}z\big)$ on the $\mathrm{Log}z$-plane $\mathbb{C}_{\mathrm{Log}z}$. So the range of virtual period map $z^{\tau}$ means the range of this lifted map. Thus, we obtain the diagram
@@
The naive “inversion problem” in this setting is to give a map from $\mathbb{C}_{w}^{\times}$ to $\mathbb{C}_{z}^{\times}$ making the diagram commutative. However, there does not exists a univalent holomorphic map which makes the diagram commutative, since the analytic continuation of any local inversion map becomes automatically the multi-valued function $z=w^{1/\tau}$ on $\mathbb{C}_{w}^{\times}$. That is, a naive solution of the classical inversion problem for the virtual period map $z^{\tau}$ by a univalent map from the range to the domain does not exist.
@@
We saw that the virtual period map $z^{\tau}$ is presented by the univalent function $w=\exp\big(\tau\mathrm{Log}z\big)$ defined on the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$ to the range $\mathbb{C}_{w}$. Then, we immediately observe that its range $\mathbb{C}_{w}$ is the quotient of the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$ by the translation group action
@@ 5. Winding of the domain $\mathbb{C}_{z}^{\times}$ of $z^{\tau}$ to the elliptic curve $E_{\tau}$
$(1)\qquad\qquad\qquad\qquad\quad\quad\quad(w;q)_{\infty}\ =\ \prod_{m=0}^{\infty}(1-wq^{m})\qquad\qquad\qquad\qquad\qquad\qquad\qquad\vskip-2.84544pt$
called $q$-shifted factorial function ([Sl] p.88, [GR] p.351). Let us show that the infinite product (1) converges absolutely and compact uniformly to a holomorphic function on the winding quotient space $\mathbb{C}_{w}\times\mathbb{D}_{q}$, whose zero-loci is the union of the divisor $\cup_{n=0}^{\infty}\{1-wq^{n}=0\}$. We shall denote the convergent function again by $(w,q)_{\infty}$.
@@
For any compact set $K\subset\mathbb{C}_{w}\times\mathbb{D}_{q}$, set $c:=\sup\{|w|\mid(w,q)\in K\}$ and $d\!:=\!\sup\{|q|\!\mid\!(w,q)\!\in\!K\}$ so that $\sup\{|wq^{n}|\!\mid\!(w,q)\!\in\!K\}\!\leq\!cd^{n}$. Since $0\!\leq\!d\!<\!1$, the bound is sumable, i.e. $\sum_{n=0}^{\infty}cd^{n}\!=\!c/(1\!-\!d)$. This implies the absolute uniform convergence of (1) to a holomorphic function on $K$ (see [WW] p.30 for infinite products). ∎
@@
$\begin{array}[]{ccccl}\!\qquad\quad\mathcal{U}_{+}(w,q)&:=&(w;q)_{\infty}&\!=\!&\prod_{m=0}^{\infty}(1-wq^{m}).\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\\ \!(2)\quad\quad\!\mathcal{U}_{-}(w,q)&:=&(w^{-1};q)_{\infty}/(1-w^{-1})&\!=\!&\prod_{m=1}^{\infty}(1-w^{-1}q^{m}).\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\\ \!\qquad\quad\mathcal{U}(w,q)&:=\!&\!\!\mathcal{U}_{+}(w,q)\ \mathcal{U}_{-}(w,q)&\!=\!&\prod_{m=0}^{\infty}(1\!-\!wq^{m})\prod_{m=1}^{\infty}(1\!-\!w^{-1}q^{m}).\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\end{array}$
Since they converge compact uniformly to holomorphic functions on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$, we can derivate them by $w$ term-wisely. Thus we obtain three infinite sums
@@
$(4)\qquad\quad\mathcal{V}_{-}(w,q)=-\mathcal{V}_{+}(w^{-1},q)-\frac{1}{1-w}\quad\text{and}\quad\mathcal{V}(w,q)=\mathcal{V}_{+}(w,q)+\mathcal{V}_{-}(w,q).\qquad\vskip-2.27626pt$
@@
$(5)\quad\qquad\qquad\mathcal{U}(w,q)/\mathcal{U}(w^{-1},q)=-w\quad\quad\mathcal{V}(w,q)\ +\ \mathcal{V}(w^{-1},q)\ +\ 1\ =\ 0.\quad\quad\quad\quad\quad$
@@
$\vskip-5.69046pt(7)\qquad\qquad\qquad\qquad\ \ \mathcal{W}(w,q)\ :=\ w\frac{\partial}{\partial w}\mathcal{V}(w,q)\ =\ \sum_{m\in\mathbb{Z}}\frac{wq^{m}}{(1-wq^{m})^{2}}.\qquad\qquad\qquad\ \$
@@
$(8)\qquad\qquad\qquad\quad\ \ \mathcal{W}(qw,q)\ =\ \mathcal{W}(w,q)\qquad\text{and}\qquad\mathcal{W}(w,q)\ =\ \mathcal{W}(w^{-1},q).\quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
@@ 1. The function $\mathcal{U}(w,q)$ gives the double product part of Jacobi triple product:
$\qquad\qquad\sum_{n=0}^{\infty}\frac{w^{n}q^{n}}{1-q^{n}}\ =\sum_{n=0}^{\infty}\frac{wq^{n}}{1-wq^{n}}\Big|_{\mathbb{D}_{w}\times\mathbb{D}_{q}}\ =\ \mathcal{V}(w,q)\Big|_{\mathbb{D}_{w}\times\mathbb{D}_{q}}\vskip-5.69046pt$
@@ 8. Appendix
$(9)\qquad\qquad\qquad\qquad\qquad\wp(u,\tau)\ =\ (2\pi\sqrt{-1})^{2}\ \mathcal{W}(w,q)\ -\ \frac{\pi^{2}}{3}E_{2}(\tau)\quad\quad\quad\quad\quad\qquad\qquad\qquad\qquad\quad$
@@
$(10)\qquad\qquad\qquad\qquad\quad\ w=\exp(2\pi\sqrt{-1}u),\qquad q=\exp(2\pi\sqrt{-1}\tau),\qquad\qquad\qquad\quad$
@@
$(11)\qquad\qquad\quad\!\zeta(u,\tau)\ =\ -2\pi\sqrt{-1}\ \mathcal{V}(w,q)\ +\ \frac{\pi^{2}}{3}E_{2}(\tau)u\ -\ \pi\sqrt{-1}\qquad\qquad\quad\quad\quad\ \ \$
$(12)\qquad\quad\quad\sigma(u,\tau)=-\frac{1}{2\pi\sqrt{-1}}\frac{\mathcal{U}(w,q)}{\eta(\tau)^{2}}\exp\Big(\frac{\pi^{2}}{6}E_{2}(\tau)u^{2}-\pi\sqrt{-1}u+\frac{\pi\sqrt{-1}}{6}\tau\Big).\quad\quad\quad$
where $\eta(\tau)$ is Dedekind eta-function.
@@
so that $G(\tau)=-\frac{\exp(\frac{\pi\sqrt{-1}}{6}\tau)}{2\pi\sqrt{-1}\eta(\tau)^{2}}$. So, we get the formula (12). ∎

### v6

@@ Abstract.
We illustrate a rank 1 model of virtual period maps and their associated winding quotients. The winding quotient is a new notion that appears in the recent study of virtual period maps due to the appearance of exponents which are imaginary numbers. It requires a reformulation of the classical inversion problem for the classical period maps. We answer to the new inversion problem by introducing a $q$-multiplicatively periodic function, whose pull-back to the winding covering space is the Weierstrass $\wp$-function up to a correction by Eisenstein series $E_{2}$. This function appears as the propagator on elliptic curves in mathematical physics.
@@ 1. Introduction
The classical theory of elliptic integral of the first kind for a Weierstrass family of elliptic curves $X\to S$ (e.g. [Si] Chap.1) defines a period map, which is a multi-valent map from the complement of the discriminant divisor $S\setminus D$ of the base space $S$ of the family to its range domain (so called, period domain). The range is isomorphic to the monodromy covering of $S\setminus D$, i.e. the quotient space of the universal covering $(S\setminus D)^{\sim}$ divided by the kernel of the monodromy representation of $\pi_{1}(S\setminus D,*)$. Then, the Jacobi’s inversion problem for the period map ([ibid] 1.13.) is to ask for a description of the inverse map from the range to the defining domain $S\setminus D$, and its answer is given by Eisenstein series (see [ibid] Section13)
@@
Actually, the pull-back of the function $\mathcal{W}(w,q)$ to the winding covering space has been studied in quite different contexts in mathematical physics either as the propagator $P(z)$ by Dijkgraaf [D] (4.44) on elliptic curves and the propagator $P_{0}^{\infty}(w_{1},w_{2};\tau,\overline{\tau})$ of BCOV theory by Si Li ([L1] Lemma 3.1, [L2] section 4 and [LZ] 3.3), or as the genus-one Bergman kernel function $B(z_{1},z_{2})$ in the topological recursion by Eynard [Ey] p.339 Examples (c.f. also Tyurin [Ty] p.185 Example. $\omega_{\alpha}$ and Fay [F] (46)). We do not know yet whether such connection of the winding phenomenon with mathmatical physics is common for non-geometric virtual period maps in general or it remains only in these test examples. We need to test further cases [S3, G-S].
@@ 2. The virtual period map $z^{\tau}$ on $\mathbb{C}_{z}^{\times}$
from the punctured $z$-plane $\mathbb{C}_{z}^{\times}$ as the domain of the map to the punctured plane $\mathbb{C}_{w}^{\times}$ as the range of the map. Obviously it is not a single-valued map. So we define the map geometrically as follows. We first choose $1\in\mathbb{C}_{z}^{\times}$ as the base point in the domain, and define the value of the map $z^{\tau}$ at $1\in\mathbb{C}_{z}^{\times}$ to be 1. Then, the values of the map are defined by analytic continuations of $\exp\!\big(\tau\log(z)\big)$ along paths on the $z$-plane $\mathbb{C}_{z}^{\times}$ starting at 1. The monodromy of this map by the counter clock-wise turn around 0 of the $z$-variable is given by the multiplication of $\exp{\big(2\pi\sqrt{-1}\tau\big)}$ on $z^{\tau}$, i.e. its multiplication in the $w$-plane $\mathbb{C}_{w}^{\times}$. Since $|\exp{\big(2\pi\sqrt{-1}\tau\big)}|\!<\!1$, the monodromy group $\big(\exp{(2\pi\sqrt{-1}\tau)}\big)^{\mathbb{Z}}$ is an infinite cyclic group and gives a faithful representation of the fundamental group $\pi_{1}(\mathbb{C}_{z}^{\times},1)\simeq\mathbb{Z}$. Then the monodromy covering space of the defining domain of the map $z^{\tau}$ is the universal covering space $\mathbb{C}_{\mathrm{Log}z}$ of $\mathbb{C}_{z}^{\times}$, where we denote by $\mathrm{Log}z$ the coordinate of the universal covering space such that $\exp(\mathrm{Log}z)=z$. On the space $\mathbb{C}_{\rm{Log}z}$, we have the covering transformation group action
@@ 3. Lifting of the map $z^{\tau}$ to the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$
Recall that $\mathrm{Log}z$ is the coordinate of the monodromy covering plane of $\mathbb{C}_{z}^{\times}$. Then, the virtual period map $z^{\tau}$ is lifted as the single-valued map $w=\exp\big(\tau\mathrm{Log}z\big)$ on the $\mathrm{Log}z$-plane $\mathbb{C}_{\mathrm{Log}z}$. So the range of virtual period map $z^{\tau}$ means the range of this lifted map. Thus, we obtain the diagram
@@
The naive “inversion problem” in this setting is to give a map from $\mathbb{C}_{w}^{\times}$ to $\mathbb{C}_{z}^{\times}$ making the diagram commutative. However, there does not exists a single-valued holomorphic map which makes the diagram commutative, since the analytic continuation of any local inversion map becomes automatically the multi-valued function $z=w^{1/\tau}$ on $\mathbb{C}_{w}^{\times}$. That is, a naive solution of the classical inversion problem for the virtual period map $z^{\tau}$ by a single-valued map from the range to the domain does not exist.
@@
We saw that the virtual period map $z^{\tau}$ is presented by the single-valued function $w=\exp\big(\tau\mathrm{Log}z\big)$ defined on the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$ to the range $\mathbb{C}_{w}$. Then, we immediately observe that its range $\mathbb{C}_{w}$ is the quotient of the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$ by the translation group action
@@ 5. Winding of the domain $\mathbb{C}_{z}^{\times}$ of $z^{\tau}$ to the elliptic curve $E_{\tau}$
$(1)\qquad\qquad\qquad\qquad\qquad\qquad\quad\quad(w;q)_{\infty}\ =\ \prod_{m=0}^{\infty}(1-wq^{m})\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \ \ \vskip-2.84544pt$
called $q$-shifted factorial function ([Sl] p.88, [GR] p.351). Let us show that the infinite product (1) converges absolutely and compact uniformly to a holomorphic function on the winding quotient space $\mathbb{C}_{w}\times\mathbb{D}_{q}$, whose zero-loci is the union of the divisor $\cup_{n=0}^{\infty}\{1-wq^{n}=0\}$. We shall denote the convergent function by the same $(w,q)_{\infty}$.
@@
For any compact set $K\subset\mathbb{C}_{w}\times\mathbb{D}_{q}$, set $c:=\sup\{|w|\mid(w,q)\in K\}$ and $d\!:=\!\sup\{|q|\!\mid\!(w,q)\!\in\!K\}$ so that $\sup\{|wq^{n}|\!\mid\!(w,q)\!\in\!K\}\!\leq\!cd^{n}$. Since $0\!\leq\!d\!<\!1$, this bound is sumable, i.e. $\sum_{n=0}^{\infty}cd^{n}\!=\!c/(1\!-\!d)<\infty$. This implies the absolute uniform convergence of (1) to a holomorphic function on $K$ (see [WW] p.30 for infinite products). ∎
@@
$\begin{array}[]{ccccl}\!\qquad\qquad\mathcal{U}_{+}(w,q)&:=&(w;q)_{\infty}&\!=\!&\prod_{m=0}^{\infty}(1-wq^{m}).\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\\ \!(2)\qquad\quad\!\mathcal{U}_{-}(w,q)&:=&(w^{-1};q)_{\infty}/(1-w^{-1})&\!=\!&\prod_{m=1}^{\infty}(1-w^{-1}q^{m}).\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\\ \!\qquad\qquad\mathcal{U}(w,q)&:=\!&\!\!\mathcal{U}_{+}(w,q)\ \mathcal{U}_{-}(w,q)&\!=\!&\prod_{m=0}^{\infty}(1\!-\!wq^{m})\prod_{m=1}^{\infty}(1\!-\!w^{-1}q^{m}).\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\end{array}$
Since they converge compact uniformly to holomorphic functions on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$, we can derivate them by the variable $w$ term-wisely. Thus we obtain three infinite sums
@@
$(4)\qquad\qquad\quad\mathcal{V}_{-}(w,q)=-\mathcal{V}_{+}(w^{-1},q)-\frac{1}{1-w}\quad\text{and}\quad\mathcal{V}(w,q)=\mathcal{V}_{+}(w,q)+\mathcal{V}_{-}(w,q).\qquad\qquad\vskip-2.27626pt$
@@
$(5)\quad\qquad\qquad\qquad\mathcal{U}(w,q)/\mathcal{U}(w^{-1},q)=-w\quad\quad\mathcal{V}(w,q)\ +\ \mathcal{V}(w^{-1},q)\ +\ 1\ =\ 0.\quad\quad\quad\quad\quad\qquad$
@@
$\vskip-5.69046pt(7)\qquad\qquad\qquad\qquad\qquad\qquad\mathcal{W}(w,q)\ :=\ w\frac{\partial}{\partial w}\mathcal{V}(w,q)\ =\ \sum_{m\in\mathbb{Z}}\frac{wq^{m}}{(1-wq^{m})^{2}}.\qquad\qquad\qquad\qquad\qquad$
@@
$(8)\qquad\qquad\qquad\qquad\mathcal{W}(qw,q)\ =\ \mathcal{W}(w,q)\qquad\text{and}\qquad\mathcal{W}(w,q)\ =\ \mathcal{W}(w^{-1},q).\quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
@@ 1. The function $\mathcal{U}(w,q)$ gives the double product part of Jacobi triple product:
$\quad\quad\sum_{n=0}^{\infty}\frac{w^{n}q^{n}}{1-q^{n}}\ =\sum_{n=0}^{\infty}\frac{wq^{n}}{1-wq^{n}}\Big|_{\mathbb{D}_{w}\times\mathbb{D}_{q}}\ =\ \mathcal{V}(w,q)\Big|_{\mathbb{D}_{w}\times\mathbb{D}_{q}}\vskip-5.69046pt$
@@ 8. Appendix
$(9)\qquad\qquad\qquad\qquad\qquad\ \ \wp(u,\tau)\ =\ (2\pi\sqrt{-1})^{2}\ \mathcal{W}(w,q)\ -\ \frac{\pi^{2}}{3}E_{2}(\tau)\quad\quad\quad\quad\quad\qquad\qquad\qquad\qquad\qquad$
@@
$(10)\qquad\qquad\qquad\qquad\qquad\qquad\ w=\exp(2\pi\sqrt{-1}u),\qquad q=\exp(2\pi\sqrt{-1}\tau),\qquad\qquad\qquad\qquad\qquad$
@@
$(11)\qquad\qquad\qquad\qquad\zeta(u,\tau)\ =\ -2\pi\sqrt{-1}\ \mathcal{V}(w,q)\ +\ \frac{\pi^{2}}{3}E_{2}(\tau)u\ -\ \pi\sqrt{-1}\qquad\qquad\qquad\quad\qquad\quad\ \ \$
$(12)\qquad\qquad\qquad\qquad\sigma(u,\tau)\ =-\frac{1}{2\pi\sqrt{-1}}\frac{\mathcal{U}(w,q)}{\eta(\tau)^{2}}\exp\Big(\frac{\pi^{2}}{6}E_{2}(\tau)u^{2}-\pi\sqrt{-1}u+\frac{\pi\sqrt{-1}}{6}\tau\Big).\quad\quad\qquad\qquad$
where $\eta(\tau)$ is Dedekind eta-function: $\exp(\frac{\pi\sqrt{-1}}{12}\tau)\prod_{m=1}^{\infty}(1-q^{m})$.
@@
so that $G(\tau)=-\frac{\exp(\frac{\pi\sqrt{-1}}{6}\tau)}{2\pi\sqrt{-1}\eta(\tau)^{2}}$. So, we get the formula (12).